# Frustration-free free fermions and beyond

### v1

@@ I Introduction
Frustration-free gapless systems exhibit features substantially different from those of frustrated systems [9, 10, 11, 12]. Recently, it has been conjectured that when a frustration-free gapless Hamiltonian is translationally invariant, the dispersion relation of low-energy excitations above the ground state is generally quadratic or softer [11, 13]. Remarkably, such quadratic dispersions allow SSB of continuous symmetries at zero temperature even in one dimension [11], which is typically forbidden by the Hohenberg-Mermin-Wagner theorem [14, 15, 16]. Also, when the ground state correlation function shows a power-law decay, the finite-size gap of gapless excitations scales as $O((\log L)^{2}/L^{2})$ [17] [^1] Under the open boundary condition, a tighter bound $O((L^{-2})$ is derived without assuming power-law decay of correlation functions [47, 48, 49, 50, 51].. To date, these intriguing properties have been studied primarily in frustration-free spin systems. It is therefore natural to ask whether fermionic frustration-free systems have similar characteristics. However, despite some studies on frustration-free fermionic systems [19, 20, 21, 22], their properties remain less comprehensively explored compared to those of spin systems.
@@
In this work, we study frustration-free fermionic systems. First, we focus on frustration-free free fermionic systems, where Hamiltonians are expressed as quadratic forms of fermionic creation and annihilation operators. We derive a necessary and sufficient condition for free fermionic systems to be frustration-free. We find a deep connection to tight-binding models with flat bands [23, 24, 25, 26, 27, 28, 29, 5]. Leveraging this condition and band theory, we prove that for gapless systems with translation symmetry, the excitation energy disperses as $O(|\bm{k}-\bm{k}_{0}|^{2})$ around some $\bm{k}_{0}$. We then extend our results to more general fermionic systems, including cases with interactions and without translation symmetry, by generalizing the Gosset-Huang inequality [30] to fermionic cases. With this extension, we show that even in interacting and/or non-translationally invariant fermionic systems, the finite-size gap of gapless excitations scales as $O((\log L)^{2}/L^{2})$, provided the ground state correlation function exhibits a power-law decay.
@@
The remainder of this paper is organized as follows. In Sec. II, we derive a necessary and sufficient condition for frustration-freeness in free fermionic systems. We also show that this condition implies that the low-energy excitation is quadratic or softer for frustration-free free fermionic systems with translation symmetry. In Sec. III, we present several illustrative examples. In Sec. IV, we discuss the case of $\text{U}(1)$-symmetry breaking. In Sec. V, we discuss the Gosset–Huang inequality for general fermionic systems.
@@ II.1 Setting
The general expressions in Eqs. (1) and (2) do not assume the translation invariance. We will restrict ourselves to translation invariant case in Sec. II.3. We also discuss the $\text{U}(1)$ broken case in Sec. IV and interacting case and in Sec. V.
@@ II.2 Condition for frustration-freeness
To investigate the possibility of more nontrivial ground states, let $\mu_{\bm{R}\alpha}>0$ ($\alpha=1,2,\cdots,A_{\bm{R}}$) be positive eigenvalues of $h_{\bm{R}}$ and $\bm{\psi}_{\bm{R}\alpha}$ be the corresponding orthonormalized eigenvectors. Also, let $-\nu_{\bm{R}\beta}<0$ ($\beta=1,2,\cdots,B_{\bm{R}}$) be negative eigenvalues and $\bm{\phi}_{\bm{R}\beta}$ be the corresponding orthonormalized eigenvectors. There might be eigenvectors with precisely zero eigenvalues but they will not be used in the following discussions. It follows by definition that
@@
Note that both $\hat{\psi}_{\bm{R}\alpha}^{\dagger}\hat{\psi}_{\bm{R}\alpha}$ and $\hat{\phi}_{\bm{R}\beta}\hat{\phi}_{\bm{R}\beta}^{\dagger}=1-\hat{\phi}_{\bm{R}\beta}^{\dagger}\hat{\phi}_{\bm{R}\beta}$ are projectors, whose eigenvalues are either 00 or $1$. Therefore, $\hat{H}_{\bm{R}}$ is positive semi-definite as long as $\mu_{\bm{R}\alpha}>0$ and $\nu_{\bm{R}\beta}>0$. Also, the ground state energy of the local Hamiltonian $\hat{H}_{\bm{R}}$ is 00, since, for each $\bm{R}$, the state $|\Phi_{\bm{R}}\rangle\coloneqq\prod_{\beta=1}^{B_{\bm{R}}}\hat{\phi}_{\bm{R}\beta}^{\dagger}|0\rangle$ satisfies $\hat{\psi}_{\bm{R}\alpha}|\Phi_{\bm{R}}\rangle=\hat{\phi}_{\bm{R}\beta}^{\dagger}|\Phi_{\bm{R}}\rangle=0$ for all $\alpha$ and $\beta$ thanks to the anti-commutation relations in Eqs. (7) and (8). Therefore, the Hamiltonian is frustration free if and only if there exists a state $|\Phi\rangle$ such that
@@
×$\times\big{(}\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}\{\hat{\phi}_{\bm{R}^{\prime}\beta},\hat{\psi}_{\bm{R}\alpha}^{\dagger}\}\hat{\psi}_{\bm{R}\alpha}-\hat{\psi}_{\bm{R}\alpha}^{\dagger}\{\hat{\psi}_{\bm{R}\alpha},\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}\}\hat{\phi}_{\bm{R}^{\prime}\beta}\big{)}=0.$ (15)
@@ II.3 Translation invariant models
If we assume the translation symmetry, we may drop the subscript $\bm{R}$ from $N_{\bm{R}}$, $A_{\bm{R}}$, $B_{\bm{R}}$, $\mu_{\bm{R}\alpha}$, and $\nu_{\bm{R}\beta}$. The tight-binding model $\hat{H}$ has $N_{\bm{R}}=N$ bands in total. Introducing the Fourier transformation by $\hat{c}_{\bm{R}\tau}^{\dagger}=\sum_{\bm{k}}\frac{1}{\sqrt{V}}e^{-i\bm{k}\cdot\bm{R}}\hat{c}_{\bm{k}\tau}^{\dagger}$ ($\tau=1,2,\cdots,N$), we find
@@
Here, $\hat{\bm{c}}_{\bm{k}}^{\dagger}=(\hat{c}_{\bm{k}1}^{\dagger},\cdots,\hat{c}_{\bm{k}N}^{\dagger})$ is the $N$-component row vector and $H_{\bm{k}}^{(+)}$ and $H_{\bm{k}}^{(-)}$ are $N$-dimensional Hermitian matrices. The commutation relation in Eq. (15) implies $[H_{\bm{k}}^{(+)},H_{\bm{k}}^{(-)}]=0$. Therefore, the band structure of the $N$-dimensional matrix $H_{\bm{k}}\coloneqq H_{\bm{k}}^{(+)}+H_{\bm{k}}^{(-)}$ can be computed by diagonalizing $H_{\bm{k}}^{(+)}$ and $H_{\bm{k}}^{(-)}$ separately. This implies that the band dispersions of $H_{\bm{k}}^{(+)}$ and $H_{\bm{k}}^{(-)}$, by themselves, are smooth along any one-dimensional curve $\bm{k}(t)$ in the Brillouin zone (see Appendix  A).
Among the eigenvalues of $H_{\bm{k}}^{(+)}$, suppose $\omega_{\bm{k}n}^{(+)}\geq 0$ ($n=1,2,\cdots,N^{(+)}$) are not identically zero on the entire Brillouin zone. Namely, $\omega_{\bm{k}n}^{(+)}>0$ at some $\bm{k}$. Similarly, the eigenvalues $\omega_{\bm{k}m}^{(-)}\leq 0$ ($m=1,2,\cdots,N^{(-)}$) of $H_{\bm{k}}^{(-)}$ are not identically zero. As a result, other $N-N^{(+)}-N^{(-)}\geq 0$ eigenvalues are completely 00. In general, the numbers of bands are bounded as
@@
The system is gapless if and only if at least one of the bands touches the zero-energy level. For example, suppose $\omega_{\bm{k}m}^{(+)}=0$ at $\bm{k}=\bm{k}_{0}^{(+)}$. Then the analyticity of $\omega_{\bm{k}m}^{(+)}$ combined with $\omega_{\bm{k}m}^{(+)}\geq 0$ implies that the form of band dispersion around $\bm{k}=\bm{k}_{0}^{(+)}$ is quadratic or softer, i.e., $\omega_{\bm{k}m}^{(+)}=O(|\bm{k}-\bm{k}_{0}^{(+)})|^{2})$. The same argument applies to $\omega_{\bm{k}m}^{(-)}\leq 0$. Namely, when $\omega_{\bm{k}m}^{(-)}=0$ at $\bm{k}=\bm{k}_{0}^{(-)}$, then $-\omega_{\bm{k}m}^{(-)}=O(|\bm{k}-\bm{k}_{0}^{(-)})|^{2})$.
@@ II.4 Flat bands
Since the Hamiltonian contains only finite-range hoppings, $H_{\bm{k}}$ is a Laurent polynomial (i.e., including both positive and negative powers) of $e^{i\bm{k}\cdot\bm{a}_{i}}$ ($i=1,\cdots,d$), where $\bm{a}_{1},\cdots,\bm{a}_{d}$ are the primitive lattice vectors [26, 27]. Thus the eigenvector $\bm{w}_{\bm{k}\ell}$ in Eq. (20) can also be chosen as a Laurent polynomial, as long as the normalization is not required. As a consequence, the corresponding Wannier-like state in real space
@@
is compactly supported around the unit cell $\bm{R}$ [26, 27]. Note that, in general, these states are not orthonormal.
Since $\hat{w}_{\bm{R}\ell}^{\dagger}$ creates a superposition of zero energy states, $[\hat{H},\hat{w}_{\bm{R}\ell}^{\dagger}]=0$, which implies
00 0 $=\langle\Phi|\hat{w}_{\bm{R}\ell}\hat{H}_{\bm{R}^{\prime}}^{(+)}\hat{w}_{\bm{R}\ell}^{\dagger}|\Phi\rangle=\sum_{\alpha=1}^{A}\mu_{\alpha}\big{|}\{\hat{\psi}_{\bm{R}^{\prime}\alpha},\hat{w}_{\bm{R}\ell}^{\dagger}\}\big{|}^{2},$ (22)
00 0 $=\langle\Phi|\hat{w}_{\bm{R}\ell}^{\dagger}\hat{H}_{\bm{R}^{\prime}}^{(-)}\hat{w}_{\bm{R}\ell}|\Phi\rangle=\sum_{\beta=1}^{B}\nu_{\beta}\big{|}\{\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger},\hat{w}_{\bm{R}\ell}\}\big{|}^{2}$ (23)
@@
Suppose that $\hat{H}_{\bm{R}}$ is positive semidefinite and $\sum_{\bm{R}\in\Lambda}\hat{H}_{\bm{R}}$ at $U=0$ has a flat band. Let $|\Phi_{0}\rangle$ be the fully polarized state in which all zero-energy single-particle states are occupied by spin-up electrons. Then, for any $U>0$, $|\Phi_{0}\rangle$ and its SU(2)-rotated version $e^{\alpha\hat{S}^{-}}|\Phi_{0}\rangle$ ($\alpha\in\mathbb{C}$) are common zero-energy ground state among all $\hat{H}_{\bm{R}}$’s. Furthermore, these states are only ground states of $\sum_{\bm{R}\in\Lambda}\hat{H}_{\bm{R}}$ if and only if the correlation function $\langle\Phi_{0}|\hat{c}_{\bm{R}\tau\uparrow}^{\dagger}\hat{c}_{\bm{R}^{\prime}\tau^{\prime}\uparrow}|\Phi_{0}\rangle$ satisfies a connectivity condition [5].

### v2

@@ I Introduction
Frustration-free gapless systems exhibit features substantially different from those of frustrated systems [9, 10, 11, 12, 13]. Recently, it has been conjectured that when a frustration-free gapless Hamiltonian is translationally invariant, the dispersion relation of low-energy excitations above the ground state is generally quadratic or softer [12, 14]. This implies that the low-energy effective theory of gapless frustration-free systems cannot be Lorentz invariant. Remarkably, such quadratic dispersions allow SSB of continuous symmetries at zero temperature even in one dimension [12], which is typically forbidden by the Hohenberg-Mermin-Wagner theorem [15, 16, 17]. Also, when the ground state correlation function shows a power-law decay, the finite-size gap of gapless excitations scales as $O(\log L)^{2}/L^{2})$ [18] [^1] Under the open boundary condition, a tighter bound $O((L^{-2})$ is derived without assuming power-law decay of correlation functions [78, 79, 80, 81, 82].. To date, these intriguing properties have been studied primarily in frustration-free spin systems. It is therefore natural to ask whether fermionic frustration-free systems have similar characteristics. However, despite some studies on frustration-free fermionic systems [20, 21, 22, 23, 24], their properties remain less comprehensively explored compared to those of spin systems.
@@
In this work, we study frustration-free fermionic systems. First, we focus on frustration-free free fermionic systems, where Hamiltonians are expressed as quadratic forms of fermionic creation and annihilation operators. We derive a necessary and sufficient condition for free fermionic systems to be frustration-free. Such a condition has been previously investigated for 1D one-band models in class BDI [23] and 1D two-band models in class AIII [24] from a different approach. Our results are more general in that no assumptions are made regarding the dimensionality of the system, the number of bands, or the symmetry classes. We find a deep connection to tight-binding models with flat bands [25, 26, 27, 28, 29, 30, 31, 32, 5]. Leveraging this condition and band theory, we prove that for gapless systems with translation symmetry, the excitation energy disperses as $O(|\bm{k}-\bm{k}_{0}|^{2})$ around some $\bm{k}_{0}$.
We also investigate the topological properties of frustration-free free fermionic models. We establish that the valence bands in these systems always admit an (over)complete set of compactly supported Wannier-type functions, provided that the valence band energy is strictly negative. The classification of topological phases with such Wannier-type functions [28] implies that stable topological phases, including Chern insulators and $\mathbb{Z}_{2}$ topological insulators, cannot be realized in frustration-free free fermionic models. In contrast, we demonstrate via a concrete model that fragile topology [33] can be realized.
We then extend our results to more general fermionic systems, including cases with interactions and without translation symmetry, by generalizing the Gosset-Huang inequality [34] to fermionic cases. With this extension, we show that even in interacting and/or non-translationally invariant fermionic systems, the finite-size gap of gapless excitations scales as $O((\log L)^{2}/L^{2})$, provided the ground state correlation function exhibits a power-law decay.
The main results of this work are presented in conjunction with a companion letter [35]. While the companion letter includes separate discussions on a frustration-free model on the honeycomb lattice and the effects of interactions in that model, the present paper establishes the comprehensive general framework. Here, we provide rigorous mathematical derivations, extend the analysis to $\text{U}(1)$ symmetry-breaking systems, detail the topological properties, and prove the finite-size gap scaling in interacting systems.
@@
The remainder of this paper is organized as follows. In Sec. II, we derive a necessary and sufficient condition for frustration-freeness in free fermionic systems. We also show that this condition implies that the low-energy excitation is quadratic or softer for frustration-free free fermionic systems with translation symmetry. In Sec. III, we present several illustrative examples. In Sec. IV, we discuss the case of $\text{U}(1)$-symmetry breaking. In Sec. V, we examine the possible stable topology realizable within frustration-free free-fermion models. In Sec. VI, we discuss the Gosset–Huang inequality for general fermionic systems.
@@ II.1 Setting
We note that the frustration-free condition derived in this work has been extended to quasi-local Hamiltonians with interaction tails decaying faster than any power law [36, 37].
The general expressions in Eqs. (1) and (2) do not assume the translation invariance. We will restrict ourselves to translation invariant case in Sec. II.3. We also discuss the $\text{U}(1)$ broken case in Sec. IV and interacting case in Sec. VI.
@@ II.2 Condition for frustration-freeness
To investigate the possibility of more nontrivial ground states, let $\mu_{\bm{R}\alpha}>0$ ($\alpha=1,2,\cdots,A_{\bm{R}}$) be positive eigenvalues of $h_{\bm{R}}$ and $\bm{\psi}_{\bm{R}\alpha}$ be the corresponding orthonormalized eigenvectors. Also, let $\nu_{\bm{R}\beta}<0$ ($\beta=1,2,\cdots,B_{\bm{R}}$) be negative eigenvalues and $\bm{\phi}_{\bm{R}\beta}$ be the corresponding orthonormalized eigenvectors. There might be eigenvectors with precisely zero eigenvalues but they will not be used in the following discussions. It follows by definition that
@@
Note that both $\hat{\psi}_{\bm{R}\alpha}^{\dagger}\hat{\psi}_{\bm{R}\alpha}$ and $\hat{\phi}_{\bm{R}\beta}\hat{\phi}_{\bm{R}\beta}^{\dagger}=1-\hat{\phi}_{\bm{R}\beta}^{\dagger}\hat{\phi}_{\bm{R}\beta}$ are projectors, whose eigenvalues are either 0 or $1$. Therefore, $\hat{H}_{\bm{R}}$ is positive semidefinite as long as $\mu_{\bm{R}\alpha}>0$ and $\nu_{\bm{R}\beta}>0$. Also, the ground state energy of the local Hamiltonian $\hat{H}_{\bm{R}}$ is 0, since, for each $\bm{R}$, the state $|\Phi_{\bm{R}}\rangle\coloneqq\prod_{\beta=1}^{B_{\bm{R}}}\hat{\phi}_{\bm{R}\beta}^{\dagger}|0\rangle$ satisfies $\hat{\psi}_{\bm{R}\alpha}|\Phi_{\bm{R}}\rangle=\hat{\phi}_{\bm{R}\beta}^{\dagger}|\Phi_{\bm{R}}\rangle=0$ for all $\alpha$ and $\beta$ thanks to the anti-commutation relations in Eqs. (7) and (8). Therefore, the Hamiltonian is frustration free if and only if there exists a state $|\Phi\rangle$ such that
@@
$\times\big(\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}\{\hat{\phi}_{\bm{R}^{\prime}\beta},\hat{\psi}_{\bm{R}\alpha}^{\dagger}\}\hat{\psi}_{\bm{R}\alpha}-\hat{\psi}_{\bm{R}\alpha}^{\dagger}\{\hat{\psi}_{\bm{R}\alpha},\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger}\}\hat{\phi}_{\bm{R}^{\prime}\beta}\big)=0.$ (15)
@@ II.3 Translation invariant models
Consider a translation-invariant frustration-free free fermionic Hamiltonian $\hat{H}=\sum_{\bm{R}}\hat{H}^{\prime}_{\bm{R}}$. Let $\hat{T}_{\bm{R}^{\prime}}$ denote the translation operator by a vector $\bm{R}^{\prime}$. Note that the translation invariance of the total Hamiltonian $\hat{H}$ does not necessarily imply that the local terms $\hat{H}^{\prime}_{\bm{R}}$ respect the translation symmetry, i.e., $\hat{T}_{\bm{R}^{\prime}}\hat{H}^{\prime}_{\bm{R}}\hat{T}_{\bm{R}^{\prime}}^{\dagger}=\hat{H}^{\prime}_{\bm{R}+\bm{R}^{\prime}}$. However, we can always construct a frustration-free decomposition $\hat{H}=\sum_{\bm{R}}\hat{H}_{\bm{R}}$ that satisfies this condition by defining
$\hat{H}_{\bm{R}}=\frac{1}{V}\sum_{\bm{R}^{\prime}}\hat{T}_{\bm{R}^{\prime}}\hat{H}^{\prime}_{\bm{R}-\bm{R}^{\prime}}\hat{T}_{\bm{R}^{\prime}}^{\dagger},$ (16)
where $V$ is the number of unit cells  [^2] Note, however, that $\hat{H}_{\bm{R}}$ constructed this way may depend on $V$.. Consequently, we may drop the subscript $\bm{R}$ from $N_{\bm{R}}$, $A_{\bm{R}}$, $B_{\bm{R}}$, $\mu_{\bm{R}\alpha}$, and $\nu_{\bm{R}\beta}$. The tight-binding model $\hat{H}$ has $N_{\bm{R}}=N$ bands in total. Introducing the Fourier transformation by $\hat{c}_{\bm{R}\tau}^{\dagger}=\sum_{\bm{k}}\frac{1}{\sqrt{V}}e^{-i\bm{k}\cdot\bm{R}}\hat{c}_{\bm{k}\tau}^{\dagger}$ ($\tau=1,2,\cdots,N$), we find
@@
Here, $\hat{\bm{c}}_{\bm{k}}^{\dagger}=(\hat{c}_{\bm{k}1}^{\dagger},\cdots,\hat{c}_{\bm{k}N}^{\dagger})$ is the $N$-component row vector and $H_{\bm{k}}^{(+)}$ and $H_{\bm{k}}^{(-)}$ are $N$-dimensional Hermitian matrices. The commutation relation in Eq. (15) implies $H_{\bm{k}}^{(+)},H_{\bm{k}}^{(-)}]=0$. Therefore, the band structure of the $N$-dimensional matrix $H_{\bm{k}}\coloneqq H_{\bm{k}}^{(+)}+H_{\bm{k}}^{(-)}$ can be computed by diagonalizing $H_{\bm{k}}^{(+)}$ and $H_{\bm{k}}^{(-)}$ separately. This implies that the band dispersions of $H_{\bm{k}}^{(+)}$ and $H_{\bm{k}}^{(-)}$, by themselves, are smooth along any one-dimensional curve $\bm{k}(t)$ in the Brillouin zone (see Appendix A).
Among the eigenvalues of $H_{\bm{k}}^{(+)}$, suppose $\omega_{\bm{k}n}^{(+)}\geq 0$ ($n=1,2,\cdots,N^{(+)}$) are not identically zero on the entire Brillouin zone. Namely, $\omega_{\bm{k}n}^{(+)}>0$ at some $\bm{k}$. Similarly, the eigenvalues $\omega_{\bm{k}m}^{(-)}\leq 0$ ($m=1,2,\cdots,N^{(-)}$) of $H_{\bm{k}}^{(-)}$ are not identically zero. As a result, other $N-N^{(+)}-N^{(-)}\geq 0$ eigenvalues are completely 0. In general, the numbers of bands are bounded as
@@
The system is gapless if and only if at least one of the bands touches the zero-energy level. For example, suppose $\omega_{\bm{k}m}^{(+)}=0$ at $\bm{k}=\bm{k}_{0}^{(+)}$. Then the analyticity of $\omega_{\bm{k}m}^{(+)}$ combined with $\omega_{\bm{k}m}^{(+)}\geq 0$ implies that the form of band dispersion around $\bm{k}=\bm{k}_{0}^{(+)}$ is quadratic or softer, i.e., $\omega_{\bm{k}m}^{(+)}=O(|\bm{k}-\bm{k}_{0}^{(+)})|^{2})$. The same argument applies to $\omega_{\bm{k}m}^{(-)}\leq 0$. Namely, when $\omega_{\bm{k}m}^{(-)}=0$ at $\bm{k}=\bm{k}_{0}^{(-)}$, then $\omega_{\bm{k}m}^{(-)}=O(|\bm{k}-\bm{k}_{0}^{(-)})|^{2})$.
@@ II.4 Flat bands
Since the Hamiltonian contains only finite-range hoppings, $H_{\bm{k}}$ is a Laurent polynomial (i.e., including both positive and negative powers) of $e^{i\bm{k}\cdot\bm{a}_{i}}$ ($i=1,\cdots,d$), where $\bm{a}_{1},\cdots,\bm{a}_{d}$ are the primitive lattice vectors [28, 29, 30]. Thus the eigenvector $\bm{w}_{\bm{k}\ell}$ in Eq. (21) can also be chosen as a Laurent polynomial, as long as the normalization is not required. As a consequence, the corresponding Wannier-like state in real space
@@
is compactly supported around the unit cell $\bm{R}$ [28, 29, 30]. Note that, in general, these states are not orthonormal.
Since $\hat{w}_{\bm{R}\ell}^{\dagger}$ creates a superposition of zero energy states, $\hat{H},\hat{w}_{\bm{R}\ell}^{\dagger}]=0$, which implies
$0$ $=\langle\Phi|\hat{w}_{\bm{R}\ell}\hat{H}_{\bm{R}^{\prime}}^{(+)}\hat{w}_{\bm{R}\ell}^{\dagger}|\Phi\rangle=\sum_{\alpha=1}^{A}\mu_{\alpha}\big|\{\hat{\psi}_{\bm{R}^{\prime}\alpha},\hat{w}_{\bm{R}\ell}^{\dagger}\}\big|^{2},$ (23)
$0$ $=\langle\Phi|\hat{w}_{\bm{R}\ell}^{\dagger}\hat{H}_{\bm{R}^{\prime}}^{(-)}\hat{w}_{\bm{R}\ell}|\Phi\rangle=\sum_{\beta=1}^{B}\nu_{\beta}\big|\{\hat{\phi}_{\bm{R}^{\prime}\beta}^{\dagger},\hat{w}_{\bm{R}\ell}\}\big|^{2}$ (24)
@@
Suppose that $\hat{H}_{\bm{R}}$ is positive semidefinite and $\sum_{\bm{R}\in\Lambda}\hat{H}_{\bm{R}}$ at $U=0$ has a flat band. Let $|\Phi_{0}\rangle$ be the fully polarized state in which all zero-energy single-particle states are occupied by spin-up electrons. Then, for any $U>0$, $|\Phi_{0}\rangle$ and its SU(2)-rotated version $e^{\alpha\hat{S}^{-}}|\Phi_{0}\rangle$ ($\alpha\in\mathbb{C}$) are common zero-energy ground state among all $\hat{H}_{\bm{R}}$’s. Furthermore, these states are the only ground states of $\sum_{\bm{R}\in\Lambda}\hat{H}_{\bm{R}}$ if and only if the correlation function $\langle\Phi_{0}|\hat{c}_{\bm{R}\tau\uparrow}^{\dagger}\hat{c}_{\bm{R}^{\prime}\tau^{\prime}\uparrow}|\Phi_{0}\rangle$ satisfies a connectivity condition [5].
@@
II.5 Frustration-freeness in momentum space
The Fourier transforms of Eqs. (10) and (11) yield
$H^{(+)}_{\bm{k}}$ $=\sum_{\alpha}[\sqrt{\mu_{\alpha}}\bm{\psi}_{\alpha}(e^{{\mathrm{i}}\bm{k}\cdot\bm{a}_{1}},\ldots,e^{{\mathrm{i}}\bm{k}\cdot\bm{a}_{d}})]$
$\quad\quad\times[\sqrt{\mu_{\alpha}}\bm{\psi}_{\alpha}(e^{{\mathrm{i}}\bm{k}\cdot\bm{a}_{1}},\ldots,e^{{\mathrm{i}}\bm{k}\cdot\bm{a}_{d}})]^{\dagger},$ (27)
$-H^{(-)}_{\bm{k}}$ $=\sum_{\beta}[\sqrt{\nu_{\beta}}\bm{\phi}_{\beta}(e^{{\mathrm{i}}\bm{k}\cdot\bm{a}_{1}},\ldots,e^{{\mathrm{i}}\bm{k}\cdot\bm{a}_{d}})]$
$\quad\quad\times[\sqrt{\nu_{\beta}}\bm{\phi}_{\beta}(e^{{\mathrm{i}}\bm{k}\cdot\bm{a}_{1}},\ldots,e^{{\mathrm{i}}\bm{k}\cdot\bm{a}_{d}})]^{\dagger},$ (28)
where the matrices $H^{(+)}_{\bm{k}}$ and $H^{(-)}_{\bm{k}}$ are defined in Eqs. (17) and (18) and $\bm{\psi}_{\alpha}$ and $\bm{\phi}_{\beta}$ are finite-degree vector polynomials in $e^{{\mathrm{i}}\bm{k}\cdot\bm{a}_{j}}$ ($j=1,2,\ldots,d$) defined by the Fourier transforms of Eqs. (5) and (6). Note that the above equations are identities of Laurent polynomials, not just an equality on the discrete momentum points in the Brillouin zone, if the system size is sufficiently large  [^3] If the side-lengths of the system are more than twice of the maximum width of local orbials $\hat{\psi}_{\bm{R}\alpha}$ and $\hat{\phi}_{\bm{R}\beta}$, then the momentum points are dense enough to determine polynomials in Eqs. (27) and (28) uniquely.. Thus, frustration-freeness requires that positive semidefinite Laurent polynomials $H^{(+)}_{\bm{k}}$ and $H^{(-)}_{\bm{k}}$ can be expressed as sums of squares of finite-degree polynomials. Conversely, if there exist finite-degree vector polynomials $\bm{\psi}_{\alpha}$ and $\bm{\phi}_{\beta}$ such that Eqs. (27) and (28) hold, we can define local orbitals $\hat{\psi}_{\bm{R}\alpha}$ and $\hat{\phi}_{\bm{R}\beta}$ by the inverse Fourier transforms of Eqs. (5) and (6). Therefore, frustration-freeness in translation-invariant free fermionic systems is equivalent to the existence of sum-of-squares decompositions of Laurent polynomials $\pm H^{(\pm)}_{\bm{k}}$.
While finding such decompositions is not a simple task in general, it is necessary that $H^{(\pm)}_{\bm{k}}$ are at least Laurent polynomials for the system to be frustration-free. This necessary condition can be used to rule out frustration-freeness in most cases.
A natural question is whether positive semidefinite Laurent polynomials $\pm H^{(\pm)}_{\bm{k}}$ always imply the existence of sum-of-squares decompositions as in Eqs. (27) and (28) and hence frustration-freeness. Actually, this question has been extensively studied in mathematics. In one or two dimensions, any positive semidefinite Laurent polynomial can be expressed as a sum of squares of polynomials [44, 45], while in three or higher dimensions $d\geq 3$, there are counterexamples [46]. Therefore, in $d=1,2$, the necessary condition above is also sufficient for frustration-freeness, while in $d\geq 3$, there are cases where the necessary condition holds but the system is not frustration-free.